Q1) [] In how many ways can 5 boys and 4 girls be seated in a row so that those of the same gender sit together?
Q3) [] A class has 12 boys and 4 girls. If 3 students are selected at random from the class, what is the probability p that they are all boys?
Q5) [] A student is to answer 8 out of 10 questions on an exam. How many choices has he if he must answer the first 3 questions?
_______________________
Q6) [] A class contains 10 men and 20 women of which half the men and half the women have brown

Found 2 solutions by solver91311, Edwin McCravy:
Answer by solver91311(24713)



Depends on whether the order of the boys and the order of the girls matters. In other words, is Billy, Bobby, Brett, Brandon, Bozo a different arrangement than Brandon, Bozo, Billy, Brett, Bobby?

If order doesn't matter, then there are 2 arrangements. Either all 5 boys are first followed by the 4 girls, or all 4 girls are first followed by the boys.

Otherwise, the boys can be arranged in ways and the girls in ways. So if the boys come first, there are ways, and if the girls come first, there are ways. Altogether, there are ways.


John

My calculator said it, I believe it, that settles it


Answer by Edwin McCravy(20064)


Q1) [] In how many ways can 5 boys and 4 girls be seated in a row so that those of the same gender sit together?
Case 1: BBBBBGGGG

5!4! = (120)(24) = 2880

Case 2: GGGGBBBBB

Same: 2880

Answer: 2880+2880 = 5760

Q3) [] A class has 12 boys and 4 girls. If 3 students are selected at random from the class, what is the probability p that they are all boys?

 12 boys CHOOSE 3     12C3     220      11
------------------ = ------ = ------ = ----
16 people CHOOSE 3    16C3     560      28

Q5) [] A student is to answer 8 out of 10 questions on an exam. How many choices has he if he must answer the first 3 questions?
The student has to answer the other 8-3=5 out of the remaining 10-3=7

7 questions CHOOSE 5 = 7C5 = 21
Edwin